# Automated Multi-Task Learning for Joint Disease Prediction on Electronic Health Records

### v2

@@ Abstract
Electronic Health Records (EHR) have become a rich source of information with the potential to improve patient care and medical research. In recent years, machine learning models have proliferated for analyzing EHR data to predict patients’ future health conditions. Among them, some studies advocate for multi-task learning (MTL) to jointly predict multiple target diseases for improving the prediction performance over single task learning. Nevertheless, current MTL frameworks for EHR data have significant limitations due to their heavy reliance on human experts to identify task groups for joint training and design model architectures. To reduce human intervention and improve the framework design, we propose an automated approach named AutoDP, which can search for the optimal configuration of task grouping and architectures simultaneously. To tackle the vast joint search space encompassing task combinations and architectures, we employ surrogate model-based optimization, enabling us to efficiently discover the optimal solution. Experimental results on real-world EHR data demonstrate the efficacy of the proposed AutoDP framework. It achieves significant performance improvements over both hand-crafted and automated state-of-the-art methods, also maintains a feasible search cost at the same time.
@@ 3.5 Derivation
We derive the final results using the trained surrogate model. Due to the huge search space, it is still not practical to use brute force search to get the global optimum. Therefore, we propose to apply a greedy method to search for near-optimal solutions. We introduce the detailed procedure in Appendix C. The high level idea is that we first randomly initialize the configuration, and then gradually improve its multi-task gain by random mutation and greedy selection. After a reasonably large number of iterations, we are able to attain the approximate solution that can perform well as shown in Section 4.2.
@@ 4.2 Performance Evaluation
Finally, we see that the three variants of AutoDP performed better than the other methods. Among them, AutoDPμ+σ performs the best, which means the balance of exploration and exploitation is the most effective strategy for training the surrogate model. For the last method, we also report the standard deviations and p-values of statistical tests (compared to MTG+DARTS), which justifies that the improvement is significant. The runtime is approximately as the same for MTG+DARTS and AutoDP and thus this is a fair comparison.
We also include more results for individual task gains of AutoDP, please refer to Appendix B.
@@ 4.4 Ablation Study
Additionally, we use disease-based grouping (Appendix D) to first assign tasks into different groups based on their medical relevance and then employ DARTS to search for the model architecture for each group. This allows us to analyze the effectiveness of automated task grouping. By comparing disease-based grouping with the searched configurations (Appendix E), we observe that AutoDP does not strictly follow medical classifications for task grouping but achieves significant performance improvements over disease-based grouping. This indicates the necessity of using an automated search algorithm to find the optimal task grouping, which surpasses human intuition.
5 Conclusions
In this paper, we propose AutoDP, an automated multi-task learning framework for joint disease prediction on EHR data. Compared to existing work, our method largely improves the design of task grouping and model architectures by reducing human interventions. Experimental results on real-world EHR data demonstrate that the proposed framework achieves significant improvement over existing state-of-the-art methods, while maintaining a feasible search cost.
@@ Appendix A Implementation
Appendix B Individual task gains
Appendix C Greedy search
Here, we show the detailed procedure of the greedy search algorithm for deriving the final MTL configuration in Algorithm 2. Given the budget $B$, we aim at searching for $B$ samples from the search space $\mathcal{C}\times\mathcal{A}$ such that the overall gain $G$ is maximized. Specifically, we first randomly initialize the population $\mathcal{P}$ that contains $B$ pairs of task combinations and architectures. Then, at every iteration, we randomly mutate one pair $(C,A)$ from the population and see whether the overall multi-task gain will increase. If so, we update $\mathcal{P}$ accordingly. After $K_{2}$ iterations, we can obtain a near-optimal solution. In practice, we also apply multiple initial populations to avoid getting stuck on local optima. Although we only get an approximate solution, our method can already achieve significant improvements over baselines as shown in Section 4.2.
1
Input: Trained surrogate model $F(\cdot)$, $B$, $K_{2}$.
Output: Searched population $\mathcal{P}$.
3 for $v=1,2,\cdots,K_{2}$ do
4       Randomly select one pair $(C,A)$ from $\mathcal{P}$;
6       Estimate $\mathcal{P}^{\prime}$ and $\mathcal{P}$ using $F(\cdot)$;
8
9 end for
10
Algorithm 2 Greedy Search
Appendix D Disease Based Grouping
@@
Appendix E Visualization of the Searched Configurations
@@
Appendix F Limitations and future work
There are some valuable future directions based on the current version of AutoDP.
First, from the application perspective, if we aim at deploying AutoDP to real-world healthcare systems, it would be advantageous to apply it to more complex problem settings. For example, the incorporation of diverse clinical data sources beyond EHR such as claims, drugs, medical images and texts will significantly enhance the practical utility of AutoDP. Additionally, considering the dynamic nature of healthcare environments with continuously updated input data and evolving tasks, adapting the surrogate model to accommodate new data and tasks would be imperative.
Second, addressing privacy concerns within healthcare systems is a promising direction. Therefore, extending AutoDP with data processing pipelines for automatic feature engineering could offer enhanced privacy safeguards and further improve its applicability in sensitive healthcare contexts.

### v3

@@ Abstract
Electronic Health Records (EHR) have become a rich source of information with the potential to improve patient care and medical research. In recent years, machine learning models have proliferated for analyzing EHR data to predict patients’ future health conditions. Among them, some studies advocate for multi-task learning (MTL) to jointly predict multiple target diseases for improving the prediction performance over single task learning. Nevertheless, current MTL frameworks for EHR data have significant limitations due to their heavy reliance on human experts to identify task groups for joint training and design model architectures. To reduce human intervention and improve the framework design, we propose an automated approach named AutoDP, which can search for the optimal configuration of task grouping and architectures simultaneously. To tackle the vast joint search space encompassing task combinations and architectures, we employ surrogate model-based optimization, enabling us to efficiently discover the optimal solution. Experimental results on real-world EHR data demonstrate the efficacy of the proposed AutoDP framework. It achieves significant performance improvements over both hand-crafted and automated state-of-the-art methods, also maintains a feasible search cost at the same time. Source code can be found via the link: https://github.com/SH-Src/AutoDP.
@@ 20      Update $\mathbf{\Theta}$ by training the surrogate model on $\mathcal{D}$;
1
Input: Trained surrogate model $F(\cdot)$, $B$, $K_{2}$.
Output: Searched population $\mathcal{P}$.
3 for $v=1,2,\cdots,K_{2}$ do
4       Randomly select one pair $(C,A)$ from $\mathcal{P}$;
6       Estimate $\mathcal{P}^{\prime}$ and $\mathcal{P}$ using $F(\cdot)$;
8
9 end for
10
Algorithm 2 Greedy Search
@@ 3.5 Derivation
We derive the final results using the trained surrogate model. Due to the huge search space, it is still not practical to use brute force search to get the global optimum. Therefore, we propose to apply a greedy method to search for near-optimal solutions. We introduce the detailed procedure in Algorithm 2. The high level idea is that we first randomly initialize the configuration, and then gradually improve its multi-task gain by random mutation and greedy selection.
Specifically, given the budget $B$, we aim at searching for $B$ samples from the search space $\mathcal{C}\times\mathcal{A}$ such that the overall gain $G$ is maximized. We first randomly initialize the population $\mathcal{P}$ that contains $B$ pairs of task combinations and architectures. Then, at every iteration, we randomly mutate one pair $(C,A)$ from the population and see whether the overall multi-task gain will increase. If so, we update $\mathcal{P}$ accordingly. After $K_{2}$ iterations, we can obtain a near-optimal solution. In practice, we also apply multiple initial populations to avoid getting stuck on local optima. Although we only get an approximate solution, our method can already achieve significant improvements over baselines as shown in Section 4.2.
@@ 4.2 Performance Evaluation
Moreover, we see that the three variants of AutoDP performed better than the other methods. Among them, AutoDPμ+σ performs the best, which means the balance of exploration and exploitation is the most effective strategy for training the surrogate model. For the last method, we also report the standard deviations and p-values of statistical tests (compared to MTG+DARTS), which justifies that the improvement is significant. The runtime is approximately as the same for MTG+DARTS and AutoDP and thus this is a fair comparison.
@@ 4.4 Ablation Study
Additionally, we use disease-based grouping (Appendix B) to first assign tasks into different groups based on their medical relevance and then employ DARTS to search for the model architecture for each group. This allows us to analyze the effectiveness of automated task grouping. By comparing disease-based grouping with the searched configurations (Appendix C), we observe that AutoDP does not strictly follow medical classifications for task grouping but achieves significant performance improvements over disease-based grouping. This indicates the necessity of using an automated search algorithm to find the optimal task grouping, which surpasses human intuition.
5 Conclusions and Future Work
In this paper, we propose AutoDP, an automated multi-task learning framework for joint disease prediction on EHR data. Compared to existing work, our method largely improves the design of task grouping and model architectures by reducing human interventions. Experimental results on real-world EHR data demonstrate that the proposed framework achieves significant improvement over existing state-of-the-art methods, while maintaining a feasible search cost. There are also some valuable future directions based on the current version of AutoDP.
First, from the application perspective, if we aim at deploying AutoDP to real-world healthcare systems, it would be advantageous to apply it to more complex problem settings. For example, the incorporation of diverse clinical data sources beyond EHR such as claims, drugs, medical images and texts will significantly enhance the practical utility of AutoDP.
Additionally, considering the dynamic nature of healthcare environments with continuously updated input data and evolving tasks, adapting the surrogate model to accommodate new data and tasks would be imperative.
Moreover, addressing privacy concerns within healthcare systems is a promising direction. Therefore, extending AutoDP with data processing pipelines for automatic feature engineering could offer enhanced privacy safeguards and further improve its applicability in sensitive healthcare contexts.
Finally, we assume all the tasks have the same input EHR data in our problem setting, which might not always be the case in practical scenarios. Chances are that, for some diseases, there are large and well-annotated data, while for the others, there are limited data available. How we should extend the current framework to handle more heterogeneous diseases/tasks remains a challenge.
@@ Appendix A Implementation
Appendix B Disease Based Grouping
@@
Appendix C Visualization of the Searched Configurations